if i know the velocity displacement equation how to find the accerlation-displacement equation


how to use the formula a=v*dv\ds

Grade:12

1 Answers

yours katarnak Suresh
11 years ago

For 11.28, you lost me after a=v*dv/dx. Since a=v*dv/dx and v=0.18v0/x, why not derive "v" to find dv/dx and multiply the result by 0.18v0/x? That''ll give you "a".

For 11.25, what''s the actual question? The distance travelled between what time and what time?